OpenStudy (anonymous):

Solve n check -5x-23=-98

OpenStudy (anonymous):

You must first add 23 to each side -5x-23+23=-98+23 you do calulations and ill help through steps

OpenStudy (anonymous):

You need to isolate the varible

OpenStudy (anonymous):

So y wouldn't I do 5x instead of the 23?

OpenStudy (anonymous):

Its the reverse of PEMDAS

OpenStudy (anonymous):

K

OpenStudy (anonymous):

46 and 121

OpenStudy (anonymous):

No, try to look at the posetive and negative signs

OpenStudy (anonymous):

[-5x] [-23] [+23]= [-98] [+23]

OpenStudy (anonymous):

the 23s cancel out because if you give 23 and taking 23 its like it was never there in the first place 23-98=-75

OpenStudy (anonymous):

so where does the 5x go?

OpenStudy (anonymous):

-5x then -5x=-75 you have to divide -5x

OpenStudy (anonymous):

by -5

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

and when you divide a negative by a negative you get a positive

OpenStudy (anonymous):

So it is??

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Its positive 15 not negative I know its complicate now but it will get easier

OpenStudy (anonymous):

Oh...I forgot the 75 was negative